OpenStudy (anonymous):

Find the equation of the quadratic function with zeros 10 and 12 and vertex at (11, -2).

OpenStudy (anonymous):

a) y = 2x^2 - 44x + 120 b) y = 2x^2 - 44x - 240 c) y = 2x^2 - 44x + 240 d) y = -2x^2 - 44x + 240 Is it a?

OpenStudy (anonymous):

a definately not. . coz when u plz x=10 in it . . y is not=0

OpenStudy (anonymous):

then which one... I'm lost.

OpenStudy (anonymous):

c or d

OpenStudy (anonymous):

oops sorry it's c :)

OpenStudy (anonymous):

look 2(100)-44(10)+240=0 so ans=c

OpenStudy (anonymous):

Oh right, thank you!